If a boll fall from H hight and in first 3 second travel x distance and in last motion of 1 second also travel x distance  so find H.

Grade:

1 Answers

Satyam Agarwal
5 years ago
Distance travelled in first three secS= ut+1/2 gt²=0+0.5*10*3² = 45Velocity after 3 sec ,V= u +gt = 0+ 10*3 = 30Distance travelled in last sec is also x i.e 45mWhen ball hits the ground it`s velocity becomes 0 , Velocity 1 sec before hitting the ground ,S =ut+1/2gt²45=u*1 + 0.5*10*1²U = 40Now, distance travelled in the middle. ,i.e, after first 3 sec and before last 1 secV²-u² = 2gh40²- 30² = 2*10*hh = 35mTotal distance travelled = 45+45+35 =125mLet me know if the answer is incorrect. Thanks.